The ages of a father and a son add up to 55. The father's age is the sons age reversed.
Answer by richwmiller(17219)

10a+b+10b+a=55
11a+11b=55
11*(a+b)=55
a+b=5

x+y=55
x=10a+b
y=10b+a
Not enough info.
There are three possibilities
05,50
14,41
23,32